please help me solve this equation:
x-3y=0
2x+3y+3z=18
x+2y-4z=10

Answer by scott8148(6628)

subtracting the first equation from the third gives 5y-4z=10

subtracting 2 times the first equation from the second gives 9y+3z=18

(4/3)(9y+3z=18) ... 12y+4z=24 ... adding 5y-4z=10 gives 17y=34 ... y=2

x-(3)*(2)=0 ... x=6

6+2(2)-4z=10 ... -4z=0 ... z=0